For the given confidence level and values of x and n, find the following.

x=47,n=98 , confidence level 99.9%

(c) Find the margin of error. Round the answers to at least four decimal places, if necessary.
The margin of error for the given data is _?
.
Answer #1

Solution :

Given that,

Point estimate = sample proportion = = x / n = 47 / 98 = 0.4796

1 - = 0.5204

Z/2 = 3.09

Margin of error = E = Z / 2 * (( * (1 - )) / n)

= 3.09 * [(0.4796 * 0.5204) / 98]

= 0.1559

The margin of error for the given data is 0.1559